OpenStudy (anonymous):

You roll a blue die and a yellow die. What is P(the sum of the dice is at least 6 | the blue die shows a 3)? Write fractions using the slash ( / ) key. Reduce fractions to their lowest terms.

OpenStudy (anonymous):

is this conditional probability?

OpenStudy (anonymous):

blue die shows a 3 is the condition right?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

lol you think so?!?.. ok this is easy to solve, your sample space reduces to only 6 cases.. in which the blue die is 3 and the yellow die from 1 to 6.. so pick out the ones which match your condition.. and tell me if you get the answer

OpenStudy (anonymous):

would it b 3?

OpenStudy (anonymous):

probability is always less than one or at the max equal to one

OpenStudy (anonymous):

well the blue die is 3 and the only number that can b added to 3 to get six is 3

OpenStudy (anonymous):

The question is atleast six, so the sum can be six and more than that.. so how many out of a total of 6 cases match that case??

OpenStudy (anonymous):

3+3=6 3+4=7 3+5=8 is that wat u mean?

OpenStudy (anonymous):

P(the sum of the dice is at least 6 | the blue die shows a 3) is read as "the probability that the sum of the dice is at least 6 GIVEN that the blue die shows a 3.

OpenStudy (anonymous):

the cases are (1,3) (2,3).. (6,3).. total 6 cases.. now tell how many from these match your question??

myininaya (myininaya):

lets think: Blue die will always be 3 the sum of the yellow die and the blue die we want to be at least 6 (means 6 0r more) Yellow Die Blue Die 1 3 does not equal 6 or more 2 3 '''' ''' 3 3 =6 4 3 >6 5 3 >6 6 3 >9 thats all the possible rows given that the blue die is 3 so there are 6 rolls in all there and only 4 are them have sum of at least 6 so the probability is 4/6=2/3

myininaya (myininaya):

>9 6*

OpenStudy (amistre64):

You roll a blue die and a yellow die. What is P(at least 6 | the blue die shows a 3)? P(AnB)/P(B) blue 1 2 3 4 5 6 1 x y 2 x e 3 x l 4 x l 5 x o 6 x w P(B) = 6/36 blue 1 2 3 4 5 6 1 y 2 e 3 x l 4 l 5 o 6 w P(AnB) = 1/36 1/36 ----- = 1/6 if i did it right 6/36

OpenStudy (amistre64):

ack ... at least .. got it

OpenStudy (anonymous):

no amist.. 2/3 is the answer..!

OpenStudy (anonymous):

Yeah. You saw your mistake. Good explanation otherwise though.

OpenStudy (amistre64):

im getting to it :) just gotta wade thru my stupidity to get there :)

myininaya (myininaya):

go to second table you need more x's

OpenStudy (anonymous):

Ahahaha I totally relate.

myininaya (myininaya):

amistre how am i going to get 230 medals?

OpenStudy (amistre64):

blue 1 2 3 4 5 6 1 x x x x x y 2 x x x x e 3 x x x l 4 x x l 5 x o 6 w P(A) = 15/36 P(AnB) = P(A) * P(B) = 15/36 * 6/36 = 90/1296 90/1296 -------- = 90/1296 = 5/72 hmmm 216/1296

OpenStudy (amistre64):

you already have 770 medals :)

OpenStudy (amistre64):

90/216 lol :) let me try that again

OpenStudy (amistre64):

i get 5/12 but i dunno ..

myininaya (myininaya):

lol

myininaya (myininaya):

hey amistre look at my chart